OpenStudy (anonymous):

A grain silo is shown below. What is the volume of grain that could completely fill this silo rounded to the nearest whole number? Use for pi. 34,597 ft3 11,532 ft3 35,669 ft3 2,146 ft3

OpenStudy (anonymous):

use 22/7

OpenStudy (anonymous):

Its volume of the cylinder + volume of the hemisphere \[V = \pi r^{2}h + \frac{ 2 }{ 3 }\pi r^{3}\]

OpenStudy (anonymous):

h = 172 and r = 8

OpenStudy (anonymous):

So pi times radius squared times height? and then 2/3 * pi * radius cubed?

OpenStudy (anonymous):

yes boss!